OpenStudy (anonymous):

A blueprint for a house shows the height of the house 3 5/8 to be inches. The actual house is 21 1/3 feet tall. What is the unit rate in feet per inch?

OpenStudy (anonymous):

Help ME!

OpenStudy (aaliyahfairgood):

make the mixed numbers into improper fractions 5 and 3/8 = 43/8 21 and 1/3 = 64/3 then find a common denominator (24)  (129/512)/24